OpenStudy (ujjwal):

Is there a proof for \[\frac{d|x|}{dx}=\frac{x}{\sqrt x^2}\] and what do we do when x is negative?

OpenStudy (anonymous):

I cant read properly !

OpenStudy (anonymous):

Can u PLZ draw it

OpenStudy (anonymous):

i think its supposed to be this.\[\frac{d}{dx}|x|=\frac{x}{\sqrt{x^2}}\]

OpenStudy (anonymous):

its clear that the derivative of |x| is -1 if x < 0, 1 if x > 0, and doesnt exists if x = 0. Similarly, its clear that:\[\frac{x}{\sqrt{x^2}}=\frac{x}{|x|}\]which is -1 if x < 0, 1 if x > 0, and doesnt exists if x = 0. So they are the same.

OpenStudy (raden):

use derivative by implicit

hartnn (hartnn):

or you could write |x| = \(\sqrt x^2\) and then diff.

hartnn (hartnn):

by chain rule

OpenStudy (ujjwal):

Thanks, @joemath314159 and @hartnn

OpenStudy (raden):

but, if u use by implicit derivative it will work to let y=|x| = sqrt(x^2) square both side, y^2 = x^2 2y dy = 2x dx dy/dx = 2x/2y = x/y = x/|x| = x/sqrt(x^2)

OpenStudy (ujjwal):

Thanks @RadEn